# VoMP: Predicting Volumetric Mechanical Property Fields

### v1

@@ Abstract
Physical simulation relies on spatially-varying mechanical properties, often laboriously hand-crafted. VoMP is a feed-forward method trained to predict Young’s modulus($E$), Poisson’s ratio($\nu$), and density($\rho$) throughout the volume of 3D objects, in any representation that can be rendered and voxelized. VoMP aggregates per-voxel multi-view features and passes them to our trained Geometry Transformer to predict per-voxel material latent codes. These latents reside on a manifold of physically plausible materials, which we learn from a real-world dataset, guaranteeing the validity of decoded per-voxel materials. To obtain object-level training data, we propose an annotation pipeline combining knowledge from segmented 3D datasets, material databases, and a vision-language model, along with a new benchmark. Experiments show that VoMP estimates accurate volumetric properties, far outperforming prior art in accuracy and speed.
@@ 1 Introduction
Accurate physics simulation is a critical part of modern design and engineering, for example, in workflows like creating Digital Twins (virtual replicas of real systems) Grieves2017, Real-2-Sim (generating digital simulation from the real world) omniverse, and Sim-2-Real (transfering policies trained in simulation to real-world deployment) rudin2021learning. However, setting up reliable simulations remains labor-intensive, partially due to the necessity to provide accurate mechanical properties throughout the volume of every object, namely the spatially-varying Young’s Modulus($E$), Poisson’s ratio($\nu$), and density($\rho$). Common 3D capture methods kerbl3Dgaussians and 3D repositories objaverse rarely contain such annotations, forcing artists and engineers to guess or copy-paste coarse material presets in a subjective, time-consuming process. We focus on automatic prediction of these parameters, addressing important limitations of prior art.
@@ 4.2 Geometry Transformer
To transfer voxel materials back to the original representation (i.e. splat means, tets for FEM simulation, quadrature points for simulation, etc.), we use nearest neighhbour interpolation as outlined in §G.1. The per-voxel latents are passed into the decoder model of MatVAE (§˜3), which yields per-voxel material triplets, as shown in Fig.˜3.
@@ 6.1 Implementation Details
Data and Training: For material annotation, we experimentally choose Qwen 2.5 VL-72B VLM bai2023qwenvlversatilevisionlanguagemodel; bai2025qwen25vltechnicalreport. We partition our MTD and GVM datasets (§5) into 80-10-10 train, validation, and test sets. See §E for data details. For rendering we use Omniverse nvidia_omniverse_replicator and Blender blender, and for DINOv2 we use an optimized implementation nvidia_nv_dinov2_2025. During training and testing, we set the maximum number of non-empty voxels per object $L_{N}=32,768$ (sampled stochastically, §˜4.2), and sparse data structures for efficiency. See §F for more details. All experiments were performed on a machine with four 80GB A100 GPUs, where training took about $12$ hours for MatVAE and 5 days for the Transformer.
@@ 6.3 Quantitative Evaluation
(b) Mechanical Property Estimates of our method significantly outperform the baselines on all metrics. Per-voxel error rate is first computed per object, then averaged across all objects in the test set to avoid weighing some objects more. Global voxel-level normalization yields similar results, see Supplement Tb.˜2.
@@
Aerographite
Carbon Fiber
(high strength)
Carbon Fiber
(high modulus)
Carbon Nitride Diamond
@@ A.1 End-to-end Examples with Simulation
We demonstrate the process that enables creating simulation-ready, realistic assets from Gaussian Splats and meshes in Fig.˜8, demonstrating convincing simulations without any fine-tuning. For example, in Fig.˜8, we first capture a video of an object, and then train a 3D Gaussian Splatting model. Then, we pass it to VoMP which estimates the mechanical properties in a couple of seconds. We then use these properties in a simulator to produce a realistic (see [Uncaptioned image]: 0:36), greatly reducing the barrier toward constructing realistic interactive digital worlds directly from our physical reality.
@@ A.2 More Mechanical Property Prediction Results
We show qualitative results for inferring mechanical property fields in Fig.˜9 and 10. We notice from Fig.˜9 row 1, column 2 that our model can pick up small details like the stem of the orange at the top of the object, which is given a different Young’s modulus, though it only spans a few voxels (see [Uncaptioned image]: 1:38). We notice from Fig.˜9 row 2, column 2 that our model finds that the space isnide the pot should be made up of properties that fall in the range of dirt, even though the inside of the pot was not observed through external renders (see [Uncaptioned image]: 1:50). We notice from Fig.˜9 row 3, column 2 that our model can tolerate some noise in assets such as the Gaussian splat of a bowl with fruits segmented from a larger Gaussian splat (see [Uncaptioned image]: 2:01). We notice from Fig.˜10 row 1, column 1 that our model can accurately predict thing segments and thin boundaries, like for the seat of the chair (see [Uncaptioned image]: 2:10). We notice from Fig.˜9 row 3, column 2 that our models can handle complex assets and complex materials like trees and accurately handle fine details, such as understanding where all the leaves lie and giving them different material properties than wood (see [Uncaptioned image]: 2:16). We notice from Fig.˜9 row 4, column 1 that our models can handle complex volumetric materials, such as annotating the properties of wood under the flowers (see [Uncaptioned image]: 2:28). We notice from Fig.˜9 row 4, column 2 that our models can handle thin materials and identify the vein of the leaves (see [Uncaptioned image]: 2:35). For completeness, we also include our metrics normalized by total test set voxels in  Tb.˜2 and observe the same performance boost compared to prior art as with the per-object normalization as presented in the main paper Fig.˜6.
@@ A.4.1 Distribution learned by MatVAE
We measure standard metrics used for measuring the difference in the distribution learned by MatVAE and the distribution of the test set in Tb. 3, observing small errors which suggest that MatVAE learned a good approximation of the true material distribution.
@@ A.4.3 Interpolation with MatVAE
Latent Dimension 2Latent Dimension 2
$2.23\times 10^{-1}$ $37\times 10^{-1} 1.1$$14\times 10^{-1}$ $1.00\times 10^{-1}$ $6.85\times 10^{-2}$
@@
Latent Dimension 2Latent Dimension 2
$66\times 10^{-1} 1.1$$1\times 10^{-1} 1.1$$0\times 10^{-1} 1.1$$10\times 10^{-1}$ $1.08\times 10^{-1}$
@@
Latent Dimension 2Latent Dimension 2
$25\times 10^{-1} 2.2$$25\times 10^{-1}$ $1.56\times 10^{-1}$ $1.51\times 10^{-1}$ $1.49\times 10^{-1}$
@@ Appendix B Comparison with Concurrent work
We compare our data annotation process with the annotation process of Pixie le2025pixiefastgeneralizablesupervised in  Fig.˜14. Our method performs annotation from meshes while Pixie le2025pixiefastgeneralizablesupervised gets points from training a NeRF, which often produces noisy points, and the segmentation is performed based on CLIP features, which often produces noisy segmentation for difficult objects.
@@
Although we do not have access to the Pixie le2025pixiefastgeneralizablesupervised dataset, Pixie uses in-context physics examples, which include material names and ranges of mechanical property triplets in the annotation process. We analyze these in-context physics examples and compare them with real material ranges from MTD (§˜5.1) in Tb.˜4. We find that some of these in-context properties might create pleasing simulations with a particular simulator but can fall outside the range of real materials.
@@
Although Pixie le2025pixiefastgeneralizablesupervised is concurrent with us, we still compare the outputs of Pixie with VoMP on one object in Fig.˜15.
@@ Appendix C Ablations
We provide an in-depth analysis motivating our MatVAE and Geometry Transformer training scheme by ablating each component in Tb.˜5 and 6. Our ablations require changing the hyperparameters for fair comparisons; thus, for each ablation, we tune our hyperparameters within an identical compute budget.
@@
Technically, MatVAE (§˜3) is built on top of the vanilla VAE kingma2022autoencodingvariationalbayes and we can use a vanilla VAE kingma2022autoencodingvariationalbayes in its place. In Tb.˜5, we show material property reconstruction and distributional metrics between a vanilla VAE and our MatVAE. MatVAE outperforms the vanilla VAE in almost all metrics. We find that Vanilla VAE collapses to the Young’s Modulus property, giving us a low reconstruction error for Young’s Modulus but significantly higher errors for other properties.
@@
For aggregating image features (§˜4.1), we experiment with using DINOv2 oquab2024dinov2learningrobustvisual, CLIP radford2021learningtransferablevisualmodels, and RGB colors by average pooling in the voxel. The results are shown in Tb.˜6. Our model had many layers in the Geometry Transformer (§˜4.2 and F.2) initialized from a generation model xiang2025structured3dlatentsscalable. This set of ablations was trained starting from random weights, due to the absence of generation weights for these settings. We find that using DINOv2 oquab2024dinov2learningrobustvisual and CLIP radford2021learningtransferablevisualmodels without initializing the weights from TRELLIS xiang2025structured3dlatentsscalable performs slightly worse, whereas simply using RGB colors performs significantly worse.
@@
We experiment with different normalization schemes (§˜3) like $Z$-score, and not using $\log$-space trasnform for either Young’s Modulus or Density. All of these normalization schemes lead to a significant degradation in performance. Most notably, removing the logarithmic scaling for Young’s Modulus (w/o $\log(E)$) or using a simple $Z$-score severely harms prediction accuracy.
@@ D.4 Interpreting Errors for Material Property Estimation
We simulate a 140 N compressive force, which is common in robotic gripping applications, for example, the Franka Emika franka_description “Hand” end effector applies a maximum of 70 N per finger with a maximum clamping force of 140 N. We demonstrate the results from 486 simulations in this setting, all of which were run to convergence in Fig.˜17.
@@
We simulate a 120 N impact force that simulates package drop scenarios, calculated from the impact dynamics of a 1 kg package dropped from 0.6 m height with a 5 cm deformation distance. We demonstrate the results from 486 simulations in this setting, all of which were run to convergence in Fig.˜18.
@@
We simulate a 330 N force corresponding to standard tensile testing conditions employed in bench-top universal testing machines ASTM_D638_2022; ASTM_E8_E8M_2024. We demonstrate the results from 486 simulations in this setting, all of which were run to convergence in Fig.˜19.
@@
We simulate a 200 N force, which represents typical pretension in tendon-driven robotic systems, where continuum arms and wearable assistive devices maintain structural stiffness through cable tensions Sch_ffer_2024. We demonstrate the results from 486 simulations in this setting all of which were run to convergence in Fig.˜20.
@@ E.1 Annotation with Vision-Language Model
To create our training dataset (§˜5) for the Geometry Transformer, we use a Vision-Language Model (VLM) coupled with multiple other data sources like 3D assets, component-wise part segmentations, material databases (§˜5.1), visual textures, and material names to annotate our dataset. We run the VLM on every segment of every object individually. We experiment with Qwen2.5-VL 7B, Qwen2.5-VL 32B, Qwen2.5-VL 72B bai2023qwenvlversatilevisionlanguagemodel; bai2025qwen25vltechnicalreport, VL-Rethinker wang2023autorecon, SpatialRGPT cheng2024spatialrgpt, and Cosmos Nemotron lin2024vilapretrainingvisuallanguage. We experimentally choose Qwen2.5-VL 72B for data annotation. We show the system prompts and user prompts we use in Fig.˜21, 22, 23 and 24. We find the best performing system prompts with TextGrad yuksekgonul2025optimizing. We show a response from the model for one of the segments from an object in our dataset in Fig.˜25 and 26.
@@
We construct a tiny dataset consisting of complex objects that are manually annotated and compare these properties with the outputs from the VLM in Tb.˜7. We observe that the VLM, given significant additional information, performs very well in annotating complex objects. We report the commonly used metrics we list in §˜D.2 for measuring differences in mechanical properties.

### v2

@@ Abstract
Physical simulation relies on spatially-varying mechanical properties, often laboriously hand-crafted. VoMP is a feed-forward method trained to predict Young’s modulus ($E$), Poisson’s ratio ($\nu$), and density ($\rho$) throughout the volume of 3D objects, in any representation that can be rendered and voxelized. VoMP aggregates per-voxel multi-view features and passes them to our trained Geometry Transformer to predict per-voxel material latent codes. These latents reside on a space of physically plausible materials, which we learn from a real-world dataset, guaranteeing the validity of decoded per-voxel materials. To obtain object-level training data, we propose an annotation pipeline combining knowledge from segmented 3D datasets, material databases, and a vision-language model, along with a new benchmark. Experiments show that VoMP estimates accurate volumetric properties, far outperforming prior art in accuracy and speed.
@@ 1 Introduction
Accurate physics simulation is a critical part of modern design and engineering, for example, in workflows like creating Digital Twins (virtual replicas of real systems) Grieves2017, Real-2-Sim (generating digital simulation from the real world) omniverse, and Sim-2-Real (transferring policies trained in simulation to real-world deployment) rudin2021learning. However, setting up reliable simulations remains labor-intensive, partially due to the necessity to provide accurate mechanical properties throughout the volume of every object, namely the spatially-varying Young’s Modulus ($E$), Poisson’s ratio ($\nu$), and density ($\rho$). Common 3D capture methods kerbl3Dgaussians and 3D repositories objaverse rarely contain such annotations, forcing artists and engineers to guess or copy-paste coarse material presets in a subjective, time-consuming process. We focus on automatic prediction of these parameters, addressing important limitations of prior art.
@@ 4.2 Geometry Transformer
To transfer voxel materials back to the original representation (i.e. splat means, tets for FEM simulation, quadrature points for simulation, etc.), we use nearest neighbour interpolation as outlined in §G.1. The per-voxel latents are passed into the decoder model of MatVAE (§˜3), which yields per-voxel material triplets, as shown in Fig.˜3.
@@ 6.1 Implementation Details
Data and Training: For material annotation, we experimentally choose Qwen 2.5 VL-72B VLM bai2023qwenvlversatilevisionlanguagemodel; bai2025qwen25vltechnicalreport. We partition our MTD and GVM datasets (§5) into 80-10-10 train, validation, and test sets. See §E for data details. For rendering we use Omniverse nvidia_omniverse_replicator and Blender blender, and for DINOv2 we use an optimized implementation nvidia_nv_dinov2_2025. During training and testing, we set the maximum number of non-empty voxels per object $L_{N}=32768$ (sampled stochastically, §˜4.2), and sparse data structures for efficiency. See §F for more details. All experiments were performed on a machine with four 80GB A100 GPUs, where training took about $12$ hours for MatVAE and 5 days for the Transformer.
@@ 6.3 Quantitative Evaluation
We are unable to make the vegetation subset of our dataset publicly available. Thus, we compute the mechanical property estimations on the public version of the dataset in Tb.˜2 and 3. We find that our results averaged over the public dataset are highly similar to the full dataset.
@@
(b) Mechanical Property Estimates of our method significantly outperform the baselines on all metrics. Per-voxel error rate is first computed per object, then averaged across all objects in the test set to avoid weighing some objects more. Global voxel-level normalization yields similar results, see Supplement Tb.˜4.
@@
Aerographite Carbon Fiber (high strength) Carbon Fiber (high modulus) Carbon Nitride Diamond
@@ 6.4 Reconstructing and Generating Materials with MatVAE
Table 2: Mechanical Property Estimates of our method on the publicly released dataset are very close to the full dataset. Per-voxel error rate is first computed per object, then averaged across all objects in the test set to avoid weighing some objects more. Global voxel-level normalization yields similar results, see Supplement Tb.˜3.
Ours 0.3794 ($\pm$0.29) 0.0409 ($\pm$0.04) 0.0241 ($\pm$0.01) 0.0818 ($\pm$0.03) 142.7017 ($\pm$166.92) 0.0921 ($\pm$0.07)
@@ 7 Discussion
Acknowledgments
We thank Gilles Daviet for help in setting up some of the simulations. We thank Jean-Francois Lafleche for help with rendering. We thank Beau Perschall for help in using the datasets.
@@ A.1 End-to-end Examples with Simulation
We demonstrate the process that enables creating simulation-ready, realistic assets from Gaussian Splats and meshes in Fig.˜8, demonstrating convincing simulations without any fine-tuning. For example, in Fig.˜8, we first capture a video of an object, and then train a 3D Gaussian Splatting model. Then, we pass it to VoMP which estimates the mechanical properties in a couple of seconds. We then use these properties in a simulator to produce a realistic (see [Uncaptioned image]: 0:36), greatly reducing the barrier toward constructing realistic interactive digital worlds directly from our physical reality. We demonstrate a Gaussian Splat scene with multiple simulated objects each of which has properties estimated with VoMP, we then place a robot in the scene interacting with the splats in Fig.˜14. Through our experiments on Gaussian Splats we find the Gaussian Splat voxelization scheme (§6.1) we introduce empirically qualitatively accurately voxelizes complex, noisy real-world splat objects.
@@ A.2 More Mechanical Property Prediction Results
We show qualitative results for inferring mechanical property fields in Fig.˜9 and 10. We notice from Fig.˜9 row 1, column 2 that our model can pick up small details like the stem of the orange at the top of the object, which is given a different Young’s modulus, though it only spans a few voxels (see [Uncaptioned image]: 1:38). We notice from Fig.˜9 row 2, column 2 that our model finds that the space inside the pot should be made up of properties that fall in the range of dirt, even though the inside of the pot was not observed through external renders (see [Uncaptioned image]: 1:50). We notice from Fig.˜9 row 3, column 2 that our model can tolerate some noise in assets such as the Gaussian splat of a bowl with fruits segmented from a larger Gaussian splat (see [Uncaptioned image]: 2:01). We notice from Fig.˜10 row 1, column 1 that our model can accurately predict thin segments and thin boundaries, like for the seat of the chair (see [Uncaptioned image]: 2:10). We notice from Fig.˜9 row 3, column 2 that our models can handle complex assets and complex materials like trees and accurately handle fine details, such as understanding where all the leaves lie and giving them different material properties than wood (see [Uncaptioned image]: 2:16). We notice from Fig.˜9 row 4, column 1 that our models can handle complex volumetric materials, such as annotating the properties of wood under the flowers (see [Uncaptioned image]: 2:28). We notice from Fig.˜9 row 4, column 2 that our models can handle thin materials and identify the vein of the leaves (see [Uncaptioned image]: 2:35). For completeness, we also include our metrics normalized by total test set voxels in Tb.˜4 and observe the same performance boost compared to prior art as with the per-object normalization as presented in the main paper Fig.˜6.
We present dataset voxel-averaged results on the publicly-released dataset in Tb.˜3. The publicly-released dataset does not include the vegetation subset.
Table 3: Voxel Mechanical Property Estimation. Errors for predicting mechanical properties on the publicly released dataset, which does not include the vegetation subset, are very close to the full dataset. These metrics are computed by averaging across all voxels across all 3D objects in the public test set.
Ours 0.3766 ($\pm$0.39) 0.0421 ($\pm$0.05) 0.0250 ($\pm$0.01) 0.0837 ($\pm$0.03) 113.4683 ($\pm$302.19) 0.0909 ($\pm$0.14)
@@ A.4.1 Distribution learned by MatVAE
We measure standard metrics used for measuring the difference in the distribution learned by MatVAE and the distribution of the test set in Tb. 5, observing small errors which suggest that MatVAE learned a good approximation of the true material distribution.
@@ A.4.3 Interpolation with MatVAE
Latent Dimension 2Latent Dimension 2 $2.23\times 10^{-1}$ $37\times 10^{-1} 1.1$$14\times 10^{-1}$ $1.00\times 10^{-1}$ $6.85\times 10^{-2}$
@@
Latent Dimension 2Latent Dimension 2 $66\times 10^{-1} 1.1$$1\times 10^{-1} 1.1$$0\times 10^{-1} 1.1$$10\times 10^{-1}$ $1.08\times 10^{-1}$
@@
Latent Dimension 2Latent Dimension 2 $25\times 10^{-1} 2.2$$25\times 10^{-1}$ $1.56\times 10^{-1}$ $1.51\times 10^{-1}$ $1.49\times 10^{-1}$
@@
Refer to caption Refer to caption Refer to caption
time
Refer to caption Refer to caption Refer to caption
Young’s Modulus Density Poisson’s Ratio
Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
Voxelized Splats
Figure 14: Simulating a Large Gaussian Splat Scene. We demonstrate an elastodynamic simulation of a large Gaussian Splat scene with multiple objects segmented out and being assigned properties with VoMP.
@@ Appendix B Comparison with Concurrent work
We compare our data annotation process with the annotation process of Pixie le2025pixiefastgeneralizablesupervised in  Fig.˜15. Our method performs annotation from meshes while Pixie le2025pixiefastgeneralizablesupervised gets points from training a NeRF, which often produces noisy points, and the segmentation is performed based on CLIP features, which often produces noisy segmentation for difficult objects.
@@
Although we do not have access to the Pixie le2025pixiefastgeneralizablesupervised dataset, Pixie uses in-context physics examples, which include material names and ranges of mechanical property triplets in the annotation process. We analyze these in-context physics examples and compare them with real material ranges from MTD (§˜5.1) in Tb.˜6. We find that some of these in-context properties might create pleasing simulations with a particular simulator but can fall outside the range of real materials.
@@
Although Pixie le2025pixiefastgeneralizablesupervised is concurrent with us, we still compare the outputs of Pixie with VoMP on one object in Fig.˜16.
@@ Appendix C Ablations
We provide an in-depth analysis motivating our MatVAE and Geometry Transformer training scheme by ablating each component in Tb.˜7 and 8. Our ablations require changing the hyperparameters for fair comparisons; thus, for each ablation, we tune our hyperparameters within an identical compute budget.
@@
Technically, MatVAE (§˜3) is built on top of the vanilla VAE kingma2022autoencodingvariationalbayes and we can use a vanilla VAE kingma2022autoencodingvariationalbayes in its place. In Tb.˜7, we show material property reconstruction and distributional metrics between a vanilla VAE and our MatVAE. MatVAE outperforms the vanilla VAE in almost all metrics. We find that Vanilla VAE collapses to the Young’s Modulus property, giving us a low reconstruction error for Young’s Modulus but significantly higher errors for other properties.
@@
For aggregating image features (§˜4.1), we experiment with using DINOv2 oquab2024dinov2learningrobustvisual, CLIP radford2021learningtransferablevisualmodels, and RGB colors by average pooling in the voxel. The results are shown in Tb.˜8. Our model had many layers in the Geometry Transformer (§˜4.2 and F.2) initialized from a generation model xiang2025structured3dlatentsscalable. This set of ablations was trained starting from random weights, due to the absence of generation weights for these settings. We find that using DINOv2 oquab2024dinov2learningrobustvisual and CLIP radford2021learningtransferablevisualmodels without initializing the weights from TRELLIS xiang2025structured3dlatentsscalable performs slightly worse, whereas simply using RGB colors performs significantly worse.
@@
We experiment with different normalization schemes (§˜3) like $Z$-score, and not using $\log$-space transform for either Young’s Modulus or Density. All of these normalization schemes lead to a significant degradation in performance. Most notably, removing the logarithmic scaling for Young’s Modulus (w/o $\log(E)$) or using a simple $Z$-score severely harms prediction accuracy.
@@ D.4 Interpreting Errors for Material Property Estimation
We simulate a 140 N compressive force, which is common in robotic gripping applications, for example, the Franka Emika franka_description “Hand” end effector applies a maximum of 70 N per finger with a maximum clamping force of 140 N. We demonstrate the results from 486 simulations in this setting, all of which were run to convergence in Fig.˜18.
@@
We simulate a 120 N impact force that simulates package drop scenarios, calculated from the impact dynamics of a 1 kg package dropped from 0.6 m height with a 5 cm deformation distance. We demonstrate the results from 486 simulations in this setting, all of which were run to convergence in Fig.˜19.
@@
We simulate a 330 N force corresponding to standard tensile testing conditions employed in bench-top universal testing machines ASTM_D638_2022; ASTM_E8_E8M_2024. We demonstrate the results from 486 simulations in this setting, all of which were run to convergence in Fig.˜20.
@@
We simulate a 200 N force, which represents typical pretension in tendon-driven robotic systems, where continuum arms and wearable assistive devices maintain structural stiffness through cable tensions Sch_ffer_2024. We demonstrate the results from 486 simulations in this setting all of which were run to convergence in Fig.˜21.
@@ E.1 Annotation with Vision-Language Model
To create our training dataset (§˜5) for the Geometry Transformer, we use a Vision-Language Model (VLM) coupled with multiple other data sources like 3D assets, component-wise part segmentations, material databases (§˜5.1), visual textures, and material names to annotate our dataset. We run the VLM on every segment of every object individually. We experiment with Qwen2.5-VL 7B, Qwen2.5-VL 32B, Qwen2.5-VL 72B bai2023qwenvlversatilevisionlanguagemodel; bai2025qwen25vltechnicalreport, VL-Rethinker wang2023autorecon, SpatialRGPT cheng2024spatialrgpt, and Cosmos Nemotron lin2024vilapretrainingvisuallanguage. We experimentally choose Qwen2.5-VL 72B for the data annotation. We show the system prompts and the user prompts that we use in Fig.˜22, 23, 24 and 25. We find the best performing system prompts with TextGrad yuksekgonul2025optimizing. We show a response from the model for one of the segments from an object in our dataset in Fig.˜26 and 27.
@@
We construct a tiny dataset consisting of complex objects that are manually annotated and compare these properties with the annotations from the VLM (Qwen2.5-VL 72B ) in Tb.˜9. We observe that the VLM, given significant additional information as we provide, performs close to human annotator performance. We report the commonly used metrics we list in §˜D.2 for measuring differences in mechanical properties.